OpenStudy (inkyvoyd):

Solve the following system of equations by elimination: x+y=30 x=25-y

OpenStudy (inkyvoyd):

@Jesstho.-.

OpenStudy (anonymous):

55

OpenStudy (crashonce):

subs the second one into the first one... 25-y+y=30 therefore 25-y = 30-y therefore there are no real solutions

OpenStudy (inkyvoyd):

55?

OpenStudy (anonymous):

WHOOPS

OpenStudy (inkyvoyd):

thank you @CrashOnce

OpenStudy (anonymous):

Sorry I read the problem wrong

OpenStudy (crashonce):

lol

OpenStudy (inkyvoyd):

@Jesstho.-. I have more

OpenStudy (anonymous):

okay

OpenStudy (adrynicoleb):

Hedgehogs Jk Kitten